OpenStudy (ammarah):

Tell whether x and y show direct variation, inverse variation, or neither. 1. y=2x+3 2. y=x/3 3. x=3/y 4. 1/2xy=2

OpenStudy (ammarah):

@Mertsj

OpenStudy (mertsj):

find some ordered pairs. If the variation is direct, y/x will always be the same number If the variation is inverse xy will always be the same number Otherwise, it is neither

OpenStudy (mertsj):

Let us look at number 1. (1,5) (2,7) (3,9) 5/1=5 7/2=3.5 Nope. Not direct 1x5=5 2x7=14 Nope. Not inverse Neither

OpenStudy (ammarah):

can u do it another way? our teacher never said anything about ordered pairs....like he said to rewrite the equation?

OpenStudy (mertsj):

I have no way of knowing what your teacher said. You should consult your notes. If the equation can be written y = kx it is direct variation If the equation can be written y = k/x, it is inverse variation Otherwise it is neither.

OpenStudy (ammarah):

how about number two? is it inverse?

OpenStudy (ammarah):

@Mertsj

OpenStudy (ammarah):

for # three i got direct and for # 4 inverse?

OpenStudy (ammarah):

@Mertsj

OpenStudy (mertsj):

You said number 2 is inverse. Is it of the form y = k/x?

OpenStudy (ammarah):

no its x/k

OpenStudy (ammarah):

wait what is k??

OpenStudy (mertsj):

A constant.

OpenStudy (ammarah):

so its neither for #2?

OpenStudy (mertsj):

y=x/3 is the same as y =(1/3)x

OpenStudy (ammarah):

direct

OpenStudy (ammarah):

can u check #3 and 4?

OpenStudy (mertsj):

3. x=3/y Multiply both sides by y. What do you get?

OpenStudy (ammarah):

y=3/x

OpenStudy (ammarah):

inverse

OpenStudy (mertsj):

yes. What about 4?

OpenStudy (ammarah):

inverse

OpenStudy (mertsj):

yes